OpenStudy (anonymous):

help pleasee! Find the equation of the line through the point (–2, –2) that is perpendicular to the line y = 2x – 1

OpenStudy (sheraz12345):

SLope will be -1/2 and all u have to do is to apply formula y – y1 = m(x – x1)

OpenStudy (anonymous):

how you apply the formula??

OpenStudy (anonymous):

wait no, sorry perpendicular= reciprocal + flip sign y= -1/2x + b -2 = -1/2(-2) + b -2 =1 + b -3 = b y = -1/2x -3

OpenStudy (sheraz12345):

@dalia_lam u r right but the authentic way to do this is to apply point slppe form formula